# Stability analysis of very high order minimization-based and Taylor-based embedded boundary treatments of discontinuous Galerkin for hyperbolic equations

### v1

@@ 3 Shifted boundary polynomial correction
In this section, the main notation of the SB polynomial correction to impose embedded boundary conditions is recalled. Due to its simpler shape, it will be then direct to make a link with the ROD polynomial correction developed in section 4. The idea of the considered SB treatment consists in imposing boundary conditions on the computational boundary rather than the real one, while retaining the high order accuracy thanks to a truncated Taylor expansion. Let $u_{D}$ be the prescribed Dirichlet boundary condition of the advected variable on the real boundary $\bar{x}$. A modified boundary condition $v_{h}$ for the computational boundary $\tilde{x}$ can be developed starting from the following Taylor expansion:
@@
Thus, a modified boundary condition can be defined including all additional corrective terms and by imposing the real boundary condition $u(\bar{x})=u_{D}$:
@@
As shown in [ciallella2023shifted], the SB polynomial correction can be extracted naturally from equation (14), meaning that the whole set of high order correction terms can be recovered completely by simply evaluating the polynomial of the boundary cell in the point $\bar{x}=\tilde{x}+d$ and by making the difference with the solution evaluated in $\tilde{x}$. Finally the modified Dirichlet boundary condition $v_{h}$ on the computational boundary can be reformulated under the following correction:
@@
where $\mathbf{u}$ are the degrees of freedom of the boundary cell. We also introduce the modified SB basis $\bm{\varphi}^{\text{SB}}$ as:
@@ 4 ROD as a polynomial correction
The Reconstructing for Off-site Data (ROD) method is a minimization-based high order embedded boundary treatment that can be used to impose Dirichlet boundary conditions on the computational boundary. The main idea of this method consists in reconstructing the solution at the computational boundary by using a constrained minimization problem that creates a modified polynomial $v_{h}$ that is as close as possible to the internal polynomial $u_{h}$ and strictly satisfies the Dirichlet condition on the real boundary.
@@
It can be noticed that the distance $\mathcal{D}$ can be chosen in different ways as long as it stays a convex function. The simplest choice, as also followed in [santos2024very, ciallella2024very], constists in taking the Euclidean distance, meaning
@@
In the following, we show first how it is possible to write the ROD method as a polynomial correction when only one boundary constraint is imposed. Finally, we show that it is possible to obtain a similar formulation also when multiple boundary constraints are imposed on the polynomial $v_{h}$, as it was done in [ciallella2024very, santos2024very].
@@ 4.1 Euclidean norm as cost function
Remark 4.2 (Similarities with the SB method).
@@
Consistency is also respected since for body fitted boundary conditions one has $d=0$ and $\bar{x}=\tilde{x}$. It is easily shown that this brings to $\alpha^{\text{ROD-E}}(\tilde{x})=1$ and $\bm{\varphi}^{\text{ROD-E}}(\tilde{x})=\mathbf{0}$, which eventually gives $v_{h}(\tilde{x})=u_{D}$.
@@ 4.4 ROD as a polynomial correction with multiple constraints
Although the analysis performed in this work is done on one-dimensional domains, the formulation of the ROD method as a polynomial correction can be expressed in a similar way also for multi-dimensional problems, meaning where there are more constraints for each boundary cell. Indeed, when working in multiple dimensions, the boundary term is no longer a single point as in 1D, but it is the surface integral computed on the computational boundary, which is in general defined as the union of the closest mesh edges, of the starting (easily generated) mesh, to the real embedded boundary . As it was shown on recent works [santos2024very, ciallella2024very], where ROD is coupled with DG, in general a single ROD polynomial for each boundary cell is found by introducing several costraints in the ROD minimization problem. Then, this polynomial is used to reconstruct the corrected value in each quadrature point of the edge integral, performed on the computational boundary.
For multi-dimensional problem, we will refer to the ROD2D method. However, it should be noticed that the same considerations are also valid for 3D problems. In this case, the minimization problem can be written as follows. Let $u_{h}$ be the polynomial approximation of degree $p$ in the boundary element $\Omega_{b}$. A modified polynomial $v_{h}$ is defined as:
@@
In earlier works related to the development of ROD for discontinuous Galerkin [santos2024very, ciallella2024very], the following linear system is solved to obtain the coefficients $\mathbf{v}$:
@@
$\bm{\alpha}^{\text{ROD2D-E}}]^{T}=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})\Phi(\bar{\mathbf{x}}))^{-1}\quad\text{ and }\quad[\bm{\varphi}^{\text{ROD2D-E}}(\tilde{\mathbf{x}})]=\bm{\varphi}(\tilde{\mathbf{x}})-\Phi(\bar{\mathbf{x}})\bm{\alpha}^{\text{ROD2D-E}}.$ (39)
@@
$\begin{split}v_{h}(\tilde{\mathbf{x}})&=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\mathbf{v}=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\left[\mathbf{u}-\Phi(\bar{\mathbf{x}})\bm{\lambda}\right]\\ &=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\left[\mathbf{u}-\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})\Phi(\bar{\mathbf{x}}))^{-1}(\Phi^{T}(\bar{\mathbf{x}})\mathbf{u}-\mathbf{u}_{D})\right]\\ &=\left[\bm{\varphi}^{T}(\tilde{\mathbf{x}})-\bm{\varphi}^{T}(\tilde{\mathbf{x}})\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})\Phi(\bar{\mathbf{x}}))^{-1}\Phi^{T}(\bar{\mathbf{x}})\right]\mathbf{u}+\bm{\varphi}^{T}(\tilde{\mathbf{x}})\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})\Phi(\bar{\mathbf{x}}))^{-1}\mathbf{u}_{D}\end{split}$ (43)
@@
$\bm{\alpha}^{\text{ROD2D-$L^{2}$}}]^{T}=\bm{\varphi}^{T}(\tilde{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}}))^{-1}\quad\text{ and }\quad[\bm{\varphi}^{\text{ROD2D-$L^{2}$}}(\tilde{\mathbf{x}})]=\bm{\varphi}(\tilde{\mathbf{x}})-\Phi(\bar{\mathbf{x}})\bm{\alpha}^{\text{ROD2D-$L^{2}$}}.$ (47)
@@
$\begin{split}v_{h}(\tilde{\mathbf{x}})&=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\mathbf{v}=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\left[\mathbf{u}-M^{-1}\Phi(\bar{\mathbf{x}})\bm{\lambda}\right]\\ &=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\left[\mathbf{u}-M^{-1}\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}}))^{-1}(\Phi^{T}(\bar{\mathbf{x}})\mathbf{u}-\mathbf{u}_{D})\right]\\ &=\left[\bm{\varphi}^{T}(\tilde{\mathbf{x}})-\bm{\varphi}^{T}(\tilde{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}}))^{-1}\Phi^{T}(\bar{\mathbf{x}})\right]\mathbf{u}+\bm{\varphi}^{T}(\tilde{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}}))^{-1}\mathbf{u}_{D}\end{split}$ (51)
@@ 5 Stability analysis
To simplify the writing of this section, the analysis will be performed with a minimal system configuration. We consider two elements containing $p+1$ degrees of freedom each. In contrast to the Von Neumann analysis, which assumes an infinite number of cells, the presence of boundary conditions makes the analysis dependent on the finite number of cells in the domain. Nevertheless, the qualitative behavior of the results remains largely unaffected by the specific number of cells. This is also typical of other analysis involving boundary conditions that have been presented in the literature [M2AN_2015__49_1_39_0]. In the analysis carried out in [ciallella2025stability] where the SB method was studied, it has been shown that only the eigenvalues of the boundary cell are the ones affected by the embedded boundary treatment. For the analysis involving both ROD and SB embedded treatments, the reference maximum CFL number is taken as that of the internal DG scheme, computed assuming a periodic domain. For simplicity, the analysis is performed taking $\Delta x=1$.
In section 5.1, the first analysis is performed by enforcing standard periodic boundary conditions to recover the classical CFL stability constraints of DG schemes. In section 5.2, resp. section 5.4, we present the analysis of the SB polynomial correction with explicit, resp. implicit, time integration. The results of the SB correction have also been presented in [ciallella2025stability] for polynomial degree $p=1,2,3$ and are shown here for completeness and extended to polynomial degree up to 6. In section 5.3, resp. section 5.5, the analysis of the ROD polynomial correction with explicit, resp. implicit, time integration is presented.
Concerning implicit integration, only implicit Euler will be considered to show in practice the beneficial effect of implicit time integration.
5.1 Standard periodic boundary conditions and explicit time integration
The DG method, following the notation of equation (7), for a system of two cells with periodic boundary conditions imposed at the left (of the first cell $\Omega_{1}=[1/2,3/2]$) and at the right (of the second cell $\Omega_{2}=[3/2,5/2]$) reads
@@ 5.1 Standard periodic boundary conditions and explicit time integration
Therefore, the global matrices for this system are
$\mathbf{M}=\begin{bmatrix}M_{1}&0\\ 0&M_{2}\end{bmatrix},\quad\text{and}\quad\mathbf{K}=\begin{bmatrix}K^{s}_{1}-K_{1}^{R}&K_{1}^{L}\\ K_{2}^{L}&K^{s}_{2}-K_{2}^{R}\end{bmatrix}$,
where, in this simplified analysis based on uniform meshes, $M_{1}=M_{2}$ and $K^{s}_{1}=K^{s}_{2}$ and they are equal to the definitions given in equation (8). The global vector of degrees of freedom is simply $\mathbf{U}(t)=[\mathbf{u}_{1}(t)\;\mathbf{u}_{2}(t)]^{T}$. Following the notation given in equation (9), interface matrices are defined as
@@
$K^{R}_{2}=\varphi_{i}^{2}(x_{5/2})\varphi_{j}^{2}(x_{5/2}),\qquad K^{L}_{2}=\varphi_{i}^{2}(x_{3/2})\varphi_{j}^{1}(x_{3/2}).$
It can be noticed that, with periodic boundary conditions, the standard CFL restriction for DG is obtained: $\text{CFL}^{p}_{\text{max}}\approx\frac{1}{2p+1}$. The maximum amplification factor for periodic boundary conditions is presented in figure 3. In the following sections, a normalized $\text{CFL}\in[0,1]$ is considered by taking into account the reference CFL values obtained in this analysis, i.e. $\frac{\Delta t}{\Delta x}=\text{CFL}^{p}_{\text{max}}\text{CFL}$. This is done to simplify the visualization of the stability regions for the embedded boundary methods.
@@
5.2 Shifted boundary treatment and explicit time integration
In this section, the homogeneous Dirichlet boundary condition is considered to be applied on a point that does not coincide with the left interface of the first (boundary) element. Since an upwind flux is considered, there is no need to impose any boundary condition of the right interface of the second element. Recall that the analysis only includes two elements to simplify the eigenspectrum visualization.
Following the notation in section 3, the analysis will consider both situations shown in figure 2 where the boundary condition is shifted to the left interface and to the right of the first cell (i.e. $d<0$ and $d>0$, with $d\in[-\Delta x,\Delta x]$). Since the analysis is performed with $\Delta x=1$ thus the distance will be varied considering $d\in[-1,1]$. Having either $d<0$ or $d>0$ is a choice that can be taken when meshing a domain and embedding an internal boundary: $d<0$ means that the elements crossed by the embedded boundary are not considered as active elements, while $d>0$ means that the elements crossed by the embedded boundary are kept as active elements of the discretization. Herein, but also in the following sections, we study both configurations to illustrate both possibilities. Due to the upwind flux, the left interface will see the flux given by the SB polynomial correction (16):
$v_{h}(x_{1/2})=u(x_{1/2})-u(x_{1/2}+d)=\sum_{j=0}^{p}(\varphi^{1}_{j}(x_{1/2})-\varphi^{1}_{j}(x_{1/2}+d))u_{1,j},$ (53)
where a homogeneous Dirichlet condition, $u_{D}=0$, is imposed on $\bar{x}$, and $u(x_{1/2}+d)$ is the value of the solution of the boundary cell on the real boundary $\bar{x}$.
In this case, the global system is
$\begin{split}M_{1}\frac{\text{d}\mathbf{u}_{1}}{\text{d}t}&=K^{s}_{1}\mathbf{u}_{1}-(K^{R}_{1}\mathbf{u}_{1}-K^{\text{SB}}_{1}\mathbf{u}_{1}),\\ M_{2}\frac{\text{d}\mathbf{u}_{2}}{\text{d}t}&=K^{s}_{2}\mathbf{u}_{2}-(K^{R}_{2}\mathbf{u}_{2}-K^{L}_{2}\mathbf{u}_{1}),\end{split}$ (54)
where the left interface matrix that introduces the SB polynomial correction is defined as
$K^{\text{SB}}_{1}=\varphi^{1}_{i}(x_{1/2})\varphi^{\text{SB},1}_{i}(x_{1/2})=\varphi^{1}_{i}(x_{1/2})(\varphi^{1}_{j}(x_{1/2})-\varphi^{1}_{j}(x_{1/2}+d)).$
Therefore, the global matrices for this system are
$\mathbf{M}=\begin{bmatrix}M_{1}&0\\ 0&M_{2}\end{bmatrix},\quad\text{and}\quad\mathbf{K}=\begin{bmatrix}K^{s}_{1}-K_{1}^{R}+K_{1}^{\text{SB}}&0\\ K_{2}^{L}&K^{s}_{2}-K_{2}^{R}\end{bmatrix}$, (55)
where the other local matrices are defined as in section 5.1.
The stability region for explicit discontinuous Galerkin schemes with SB polynomial correction is shown in figure 4. As in the previous section, the stability region is given by the amplification factor computed from the eigenvalues of the discretized operators, which are amplified with the corresponding high order time integration method (11). It should be noticed that the stability region is not symmetric with respect to the value of $d\in[-1,1]$. When $d\in[-1,0]$, there are stable areas with a constraint on the CFL number that becomes stricter as $d\rightarrow-1$ and $p$ increases. While for $d\in[0,1]$ the stability region of the method is not constrained by the CFL number, but the method becomes unstable after a certain threshold $d>d_{max}$, which depends on the polynomial degree. In particular, when d=1d=-1, the maximum admissible time step is dictated by a $\text{CFL}\leq 0.6$ for $\mathbb{P}^{1}$, $\text{CFL}\leq 0.2$ for $\mathbb{P}^{2}$, $\text{CFL}\leq 0.05$ for $\mathbb{P}^{3}$, $\text{CFL}\leq 0.008$ for $\mathbb{P}^{4}$. For $\mathbb{P}^{5}$, the furthest distance is d=0.25d=-0.25 with an associated $\text{CFL}\leq 0.08$. For $\mathbb{P}^{6}$, the furthest distance is d=0.07d=-0.07 with an associated $\text{CFL}\leq 0.3$. This means that for $\mathbb{P}^{5}$ and $\mathbb{P}^{6}$, also for $d\in[-1,0]$, the SB method becomes unconditionally unstable after a certain distance threshold. Since, for $d\in[0,1]$, the scheme becomes unconditionally unstable after a certain distance threshold even for lower order polynomials, the configuration $d>0$ is in general not recommended.
Figure 4: Stability region: SB polynomial correction of the homogeneous Dirichlet condition for the explicit discontinuous Galerkin method. Green areas are stable, while red areas are unstable.
We can analytically compute the eigenvalues of the SB boundary discretized operator. As it was shown in [ciallella2025stability], only the eigenvalues of the boundary cell are affected by the embedded boundary condition. Therefore, we only show here the eigenvalues depending on the distance $d$. Since an analytical study of high order methods is really cumbersome, we only present here the analytical function obtained for a $\mathbb{P}^{1}$ method:
$\lambda^{\text{SB}}_{1,2}=3d-2\pm\sqrt{9d^{2}-12d-2}.$
It can be noticed that the scheme is stable in the semi-discrete sense, i.e. the real part of $\lambda_{1,2}$ is negative, if $d<2/3\approx 0.66$, which is the distance threshold after which the $\mathbb{P}^{1}$ scheme becomes unconditionally unstable.

### v2

@@ 4 Shifted boundary polynomial correction
This section briefly reviews the SB polynomial correction [8] to establish notation and provide context for the ROD formulations derived in section 5. A comprehensive stability analysis of the SB method is available in [6] and is not repeated here for brevity. The SB approach enforces high order boundary conditions on the computational boundary $\tilde{x}$ by applying a corrected Dirichlet value thanks to a truncated Taylor expansion. Let $u_{D}$ be the Dirichlet boundary condition on the real boundary $\bar{x}$, which does not coincide with $x_{L}$. A corrected boundary condition $v_{h}$ for the computational boundary $\tilde{x}$ can be constructed starting from the Taylor expansion:
@@
Hence, a modified boundary condition can be developed including all high order derivative terms and by enforcing the prescribed boundary condition $u(\bar{x})=u_{D}$:
@@
As shown in [8], from equation (12) it is direct to infer that the sum of derivative terms in the above expression is equivalent to the difference of the polynomial approximation evaluated at the two boundary locations. Therefore, the SB Dirichlet boundary condition $v_{h}$ on $\tilde{x}$ can be rewritten using the following correction:
@@
where $\mathbf{u}$ are the modal coefficients of the boundary element. Here, we also introduce the modified SB basis $\bm{\varphi}^{\text{SB}}$ as:
@@ 5 ROD as a polynomial correction
The Reconstruction for Off-site Data (ROD) method is a minimization-based high order embedded boundary treatment that can be used to impose Dirichlet boundary conditions on the computational boundary. The main idea of this method consists in reconstructing the solution at the computational boundary by using a constrained minimization problem that creates a modified polynomial $v_{h}$ that is as close as possible to the internal polynomial $u_{h}$ and strictly satisfies the Dirichlet condition on the real boundary.
@@
It can be noticed that the distance $\mathcal{D}$ can be chosen in different ways as long as it stays a convex function. The simplest choice, as also followed in [29, 7], constists in taking the Euclidean distance, meaning
@@
In the following, we show first how it is possible to write the ROD method as a polynomial correction when only one boundary constraint is imposed. Finally, we show that it is possible to obtain a similar formulation also when multiple boundary constraints are imposed on the polynomial $v_{h}$, as it was done in [7, 29].
@@ 5.1 Euclidean norm as cost function
Remark 5.2 (Similarities with the SB polynomial correction).
@@
Consistency is also maintained, as for body-fitted boundary conditions, $d=0$ and $\bar{x}=\tilde{x}$. It is easily shown that this brings to $\alpha^{\text{ROD-E}}(\tilde{x})=1$ and $\bm{\varphi}^{\text{ROD-E}}(\tilde{x})=\mathbf{0}$, which eventually gives $v_{h}(\tilde{x})=u_{D}$.
@@ 5.4 ROD as a polynomial correction with multiple constraints
Although the analysis performed in this work is done on one-dimensional domains, the formulation of the ROD method as a polynomial correction can be expressed in a similar way also for multi-dimensional problems, meaning where there are more constraints for each boundary cell. Indeed, when working in multiple dimensions, the boundary term is no longer a single point as in 1D, but it is the surface integral computed on the computational boundary, which is in general defined as the union of the closest mesh edges, of the starting (easily generated) mesh, to the real embedded boundary . As it was shown on recent works [29, 7], where ROD is coupled with DG, in general a single ROD polynomial for each boundary cell is found by introducing several costraints in the ROD minimization problem. Then, this polynomial is used to reconstruct the corrected value in each quadrature point of the edge integral, performed on the computational boundary.
The following multi-dimensional formulation is presented under the name ROD2D method, and focuses on the Euclidean and $L^{2}$ norms. However, it should be noticed that the same considerations are also valid for 3D problems. In this case, the minimization problem can be written as follows. Let $u_{h}$ be the polynomial approximation of degree $p$ in the boundary element $\Omega_{b}$. A modified polynomial $v_{h}$ is defined as:
@@
In earlier works related to the development of ROD for discontinuous Galerkin [29, 7], the following linear system is solved to obtain the coefficients $\mathbf{v}$:
@@
$\begin{split}[\bm{\alpha}^{\text{ROD2D-E}}]^{T}&=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})\Phi(\bar{\mathbf{x}}))^{-1},\\ [\bm{\varphi}^{\text{ROD2D-E}}(\tilde{\mathbf{x}})]&=\bm{\varphi}(\tilde{\mathbf{x}})-\Phi(\bar{\mathbf{x}})\bm{\alpha}^{\text{ROD2D-E}}.\end{split}$ (37)
@@
$\begin{split}v_{h}(\tilde{\mathbf{x}})&=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\mathbf{v}=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\left[\mathbf{u}-\Phi(\bar{\mathbf{x}})\bm{\lambda}\right]\\ &=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\left[\mathbf{u}-\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})\Phi(\bar{\mathbf{x}}))^{-1}(\Phi^{T}(\bar{\mathbf{x}})\mathbf{u}-\mathbf{u}_{D})\right]\\ &=\left[\bm{\varphi}^{T}(\tilde{\mathbf{x}})-\bm{\varphi}^{T}(\tilde{\mathbf{x}})\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})\Phi(\bar{\mathbf{x}}))^{-1}\Phi^{T}(\bar{\mathbf{x}})\right]\mathbf{u}+\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\bm{\varphi}^{T}(\tilde{\mathbf{x}})\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})\Phi(\bar{\mathbf{x}}))^{-1}\mathbf{u}_{D}\end{split}$ (41)
@@
$\begin{split}[\bm{\alpha}^{\text{ROD2D-$L^{2}$}}]^{T}&=\bm{\varphi}^{T}(\tilde{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}}))^{-1},\\ [\bm{\varphi}^{\text{ROD2D-$L^{2}$}}(\tilde{\mathbf{x}})]&=\bm{\varphi}(\tilde{\mathbf{x}})-\Phi(\bar{\mathbf{x}})\bm{\alpha}^{\text{ROD2D-$L^{2}$}}.\end{split}$ (45)
@@
$\begin{split}v_{h}(\tilde{\mathbf{x}})&=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\mathbf{v}=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\left[\mathbf{u}-M^{-1}\Phi(\bar{\mathbf{x}})\bm{\lambda}\right]\\ &=\bm{\varphi}^{T}(\tilde{\mathbf{x}})\left[\mathbf{u}-M^{-1}\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}}))^{-1}(\Phi^{T}(\bar{\mathbf{x}})\mathbf{u}-\mathbf{u}_{D})\right]\\ &=\left[\bm{\varphi}^{T}(\tilde{\mathbf{x}})-\bm{\varphi}^{T}(\tilde{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}}))^{-1}\Phi^{T}(\bar{\mathbf{x}})\right]\mathbf{u}+\\ &\qquad\qquad\qquad\qquad\qquad\qquad\bm{\varphi}^{T}(\tilde{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}})(\Phi^{T}(\bar{\mathbf{x}})M^{-1}\Phi(\bar{\mathbf{x}}))^{-1}\mathbf{u}_{D}\end{split}$ (49)
@@
Remark 5.5 (ROD 1D correction for multi-D problems).
In this section, we have focused on the multi-constraints ROD approach developed in previous works on ROD-DG methods [29, 7]. However, it should be noticed that by writing the 1D ROD method as the polynomial corrections (21) and (29), the ROD method in multiple dimensions could be also developed by employing several of the 1D polynomial corrections applied locally in each quadrature point. On one side, this would keep the minimization-based approach of the ROD method also in multiple dimension without the need of including multiple constraints in the reconstruction. On the other, it simplifies even more the implementation and the coding effort needed to integrate it in a general computational framework, making it as efficient as the SB polynomial correction. Future works will be dedicated to test this alternative formulations in multiple space dimensions.
@@ 6 Stability analysis
The eigenspectrum analysis is performed with a minimal system configuration. We consider two elements containing $p+1$ modal coefficients each. Due to Dirichlet boundary conditions the analysis cannot consider an infinite domain, as done for the standard Von Neumann analysis, and will be dependent on the number of elements considered in the global system. However, as it was proven in [6] for a similar configuration, the qualitative behavior of the results remains largely unaffected by the specific number of cells. This is also typical of other analysis involving boundary conditions that have been presented in the literature [32]. As a matter of fact, the analysis in [6] of the SB method showed that only the eigenvalues of the boundary cell are affected by the embedded boundary treatment. To simplify the visualization, for the analysis involving the embedded treatments, the reference maximum CFL number is taken as that of the internal DG scheme, computed assuming a periodic domain. The analysis is performed taking $\Delta x=1$.
In section 6.1, the analysis is performed by considering periodic boundary conditions, and, as expected, this allows us to recover the classical CFL stability constraints of DG schemes. In section 6.2, resp. section 6.3, the analysis of the ROD polynomial correction with explicit, resp. implicit, time integration is presented.
Implicit Euler as time integration is also studied to show whether this can have a positive impact on the stability regions of the schemes, as it was observed in [6] for the SB method.
6.1 Periodic boundary conditions and explicit time integration
The DG method, following the notation of equation (7), for a system of two elements with periodic boundary conditions imposed at the left interface of the first cell $\Omega_{1}=[1/2,3/2]$, and at the right interface of the second cell $\Omega_{2}=[3/2,5/2]$, reads
@@ 6.1 Periodic boundary conditions and explicit time integration
In this case, the global matrices are
$\mathbf{M}=\begin{bmatrix}M_{1}&0\\ 0&M_{2}\end{bmatrix},\quad\text{and}\quad\mathbf{K}=\begin{bmatrix}K^{s}_{1}-K_{1}^{R}&K_{1}^{L}\\ K_{2}^{L}&K^{s}_{2}-K_{2}^{R}\end{bmatrix}$.
Under the assumption of a uniform mesh, the local matrices are identical ($M_{1}=M_{2}$ and $K^{s}_{1}=K^{s}_{2}$) as defined in (8). The interface terms $K^{R}_{1}$ and $K^{R}_{2}$ follow from (9), and more precisely are given by
@@
$K^{R}_{2}=\varphi_{i}^{2}(x_{5/2})\varphi_{j}^{2}(x_{5/2}),\qquad K^{L}_{2}=\varphi_{i}^{2}(x_{3/2})\varphi_{j}^{1}(x_{3/2}),$
with $K^{L}_{1}$ and $K^{L}_{2}$ providing the periodic coupling between cells 1-2 and 2-1, respectively. The global vector of modal coefficients is $\mathbf{U}(t)=[\mathbf{u}_{1}(t)\;\mathbf{u}_{2}(t)]^{T}$.
As expected, when taking periodic conditions, the standard CFL condition for DG methods is obtained: $\text{CFL}^{p}_{\text{max}}\approx\frac{1}{2p+1}$. For completeness, we show in figure 2 the maximum amplification factor for this configuration. In the following sections, to simplify the visualization of stability regions, a normalized $\text{CFL}\in[0,1]$ is defined by scaling the actual CFL value with the periodic-case CFL such that $\frac{\Delta t}{\Delta x}=\text{CFL}^{p}_{\text{max}}\cdot\text{CFL}$.